OpenStudy (anonymous):

Average hourly earnings of production workers at Nussbaum Steel from 1996 through 2002 are approximated by the equation y= 0.494x- 974.5, where x represents the year and y represents the hourly earnings. What are the approximate average hourly earnings in 1996 ? (rounded to the nearest cent.)

OpenStudy (anonymous):

Solve y at 1996 and solve y at 2002. Than do this to find the average rate of change. \[\frac{ y(2002) -y(1996) }{ 2002 -1996 }\]

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

evaluate y at 2002 meaning substitute x as 2002 and solve for y. Do the same for 1996.

OpenStudy (anonymous):

Alright so what you do is this. y(2002) = 0.494x -974.5 = 0.494(2002) -(974.5) = $ 14.4 y(1996) = 0.494(1996) -974.5 = $ 11.524 hours And 2002 -1996 = 6 hours so the approximate average hourly earnings = (14.4 - 11.524)/6 = $ 0.47933/ hours Rounded to the nearest cent would be 0.48 cents.

OpenStudy (anonymous):

thx